OpenStudy (anonymous):

Estimate a 20% tip on a $31.90 bill?

OpenStudy (anonymous):

Help Please?

OpenStudy (anonymous):

😭

OpenStudy (study312):

turn 20% to a decimal

OpenStudy (study312):

can you do that?

OpenStudy (anonymous):

Is It 6.38

OpenStudy (study312):

yes by 0.20 • 31.90

OpenStudy (anonymous):

Can you help me on a couple of other questions

OpenStudy (study312):

I can double check on the tip calculator on my phone if you want. ok

OpenStudy (study312):

my phone says it is correct. what is your next question?

OpenStudy (anonymous):

One worker can pack 140 boxes in 6 hours. A second worker can pack 25% more boxes in 11 hours. Which worker packs more boxes per hour?

OpenStudy (anonymous):

worker 1

OpenStudy (study312):

show how

OpenStudy (anonymous):

well easy worker 2 takes nearly twice as long if that is the case then it should be like 80-90% faster and then it would be worker 2 but the worker 2 is only 25% faster

OpenStudy (study312):

140 ÷ 6 = 23 ( by rounding 23.3) 0.25 • 11 = 2 ( by rounding 2.4)

OpenStudy (anonymous):

well easy worker 2 takes nearly twice as long if that is the case then it should be like 80-90% faster and then it would be worker 2 but the worker 2 is only 25% faster

OpenStudy (anonymous):

medal????